A bus of wheels of diameter 1.4 if just trick kilometer then 1 minutes how much round need
Answer by Theo(13342)

the diameter of the wheels of the bus is 1.4 meters.

i'm assuming meters because that's the only measurement that makes sense.

this means the radius of the wheels of the bus is 1.4 / 2 = .7 meters.

this means the circumference of the wheels of the bus is 2 * pi * .7 = 4.398229715 meters.

one revolution of the tires will travel a linear distance of 4.398229715 meters along the ground.

if the truck travels at 60 kilometers per hour, then the bus is traveling at 1 kilometer per minute.

this is only an assumption since you did not specify how fast the bus was moving.

that is an important ingredient to this problem.

you can modify the results based on whatever speed you know that the bus was moving.

1 kilometer is equal to 1000 meters.

therefore, in one minute, the bus has traveled 1000 meters along the ground, and the wheels have rotated 1000 / 4.398229715 = 227.3642044 times.